OpenStudy (delaneyparris):

can someone check my answer plz what is a slope f a vertical line? a.-1 b.0 c.1 d.undefined***

OpenStudy (delaneyparris):

of

OpenStudy (anonymous):

ok,what's ur answer?

OpenStudy (anonymous):

what do you think?

OpenStudy (delaneyparris):

d

OpenStudy (anonymous):

Correct :)

OpenStudy (delaneyparris):

because there are no numbers so you cant really get any thing out of it. so its undefined

OpenStudy (anonymous):

right

OpenStudy (delaneyparris):

ok thanks

OpenStudy (delaneyparris):

could you help me with one more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no problem

OpenStudy (delaneyparris):

A .1/55 B. 110/1 C 1/275 D. 55/1

OpenStudy (anonymous):

what's your idea?

OpenStudy (delaneyparris):

A

OpenStudy (anonymous):

no, 165-110 = 55 220 - 165 = 55 275-220 = 55 u can see that the rate of the change is 55 and u can see that it's constant

OpenStudy (anonymous):

so it's not 1/55 . it's 55 itself

OpenStudy (delaneyparris):

so its 55/1

OpenStudy (anonymous):

yes

OpenStudy (delaneyparris):

because 55/1 is 55